How do I solve 2x^5=100 using a logarithm with an answer to three significant digits?
Answer by lwsshak3(11628)

How do I solve 2x^5=100 using a logarithm with an answer to three significant digits?
..
2x^5=100
log2+5logx=log100=2
5logx=2-log2
logx=(2-log2)/5=.33979
convert to exponential form
x=10^.33979=2.187 (the base(10) raised to the logarithm(.33979) of the number,x, is equal to the number,x(2.187)
ans:
x=2.19
..
note: you could have just as easily solved this with simple algebra.
x^5=50
raise both side to (1/5) power
(x^5)^(1/5)=50^(1/5)
x=2.19